# Identifying the impact of local connectivity patterns on dynamics in excitatory-inhibitory networks

### v2

@@ Introduction
These seminal theoretical results on excitatory-inhibitory networks however are derived via a key simplifying assumption. In standard models, the strength of the synaptic coupling between any two neurons depends on their types, but is otherwise assumed to be an independent random variable uncorrelated across synapses. This assumption is typically used to reduce a full network to a simpler circuit model that describes how the mean activities of different populations interact through averaged synaptic weights (Fig 1(a,b)). Recent synaptic-resolution experimental datasets from various species and brain areas have however revealed the prevalence of non-trivial connectivity patterns [22, 23, 24, 25, 26, 27, 28, 29, 30]. In particular, a recently released dataset from mice and humans [23, 30] reported the prominence of second-order motifs - specific correlations between pairs of synapses, such as reciprocal, chain, convergent, and divergent motifs (Fig 1(c)) - that go well beyond what is expected from independent connections, and highlights the need for a theoretical understanding of the effects of such patterns. While the presence of reciprocal connectivity motifs has been long recognized and examined within network models [29], several studies have found their influence on recurrent dynamics to be relatively modest [31, 30, 32, 33]. Other studies have argued for the importance of several types of synaptic motifs working together to determine statistical properties of network activity, such as average synchrony or correlation among neurons and network-wide dimensionality [34, 35, 32, 30, 36, 37, 38, 39, 40]. In a general theoretical analysis [35], chain motifs were found to have a dominant role in determining population-averaged responses of networks to their inputs. Nevertheless, the impact of chain motifs on excitation-inhibition balance – including central issues of stability and paradoxical responses – as well as the general interplay among these motifs and structures established by cell-type specific connectivity, remain open and intriguing questions.
@@
Figure 1: Schematic of the multi-population network model. (a) The network consists of $P$ populations, each represented by a different color. This population structure defines the statistics of synaptic connectivity. (b) The corresponding connectivity matrix consists of $P^{2}$ blocks. Synaptic weights within each block share identical statistical properties. (c) Four different types of second-order motifs, corresponding to different pair-wise correlations between synaptic weights. (d) A low-rank approximation of the connectivity matrix can integrate both the population structure and the pair-wise motif statistics.
@@
In this study, we examine how the interaction between population structure and synaptic motifs influences recurrent dynamics and the presence of paradoxical responses. To this end, we expand a previously introduced theoretical framework that allows us to reduce large networks of multiple populations to a low-dimensional description that incorporates connectivity statistics beyond the mean via a low-rank approximation [33]. Applying this theory to excitatory-inhibitory networks with chain motifs, we demonstrate that these connectivity patterns significantly impact the eigen-spectrum of the connectivity matrix, and thereby the overall recurrent dynamics. Specifically, we found that chain motifs can create strong positive feedback even in inhibition-dominated networks, and are therefore a source of a potential instability that requires revisiting the classical conditions of excitation-inhibition balance. Moreover we show that chain motifs strongly influence the responses of different populations to external inputs, and can control whether the responses are paradoxical or not. These findings highlight the intricacy of the relationship between the responses to inputs and the underlying connectivity, and in particular sound a note of caution for interpreting results of experiments in terms of only average connectivity strengths among E and I cells.
@@ I.1 Network connectivity
Following a common approach [29, 41, 34, 36, 37, 38, 39, 32], we begin by characterizing connectivity through the marginal distributions $\mathcal{P}(J_{ij}=J)$ of individual synaptic weights, and progressively incorporate higher-order correlations. In this study, we consider the first and second orders, that is, the marginal distribution of individual synaptic weights and the pair-wise correlations. We denote as motifs particular pair-wise correlations between synaptic weights (Fig 1(c)). We specifically focus on chain motifs corresponding to correlations $\tau^{c}_{ijk}$ between synapses $J_{ij}$ and $J_{jk}$ which share the intermediate neuron $j$, but $i\neq k$:
@@
Here, and throughout this study, square brackets denote the average over the full connectivity distribution.
@@
Previous studies have considered two additional types of correlations corresponding to convergent and divergent motifs. We show in Appendix. A that, in the limit of large networks, they do not contribute to the dynamics we study here.
@@ I.2 Network dynamics
Our aim is to determine how external inputs to different populations affect the steady-state activity. To this end, we will examine the response matrix $\mathbf{X}$, where the entry $\chi_{ij}$ represents the change of the steady-state activity of neuron $i$ resulting from a change in the input current to neuron $j$:
@@
$\mathbf{X}=(\mathbf{1}-\mathbf{J})^{-1}.$ (10)
The diagonal entries of the response matrix, denoted as $\chi_{ii}$, represent the activity change of neuron $i$ in response to the direct external input they receive. When this term is negative, this indicates that the neuron increases (decreases) its firing rate with a decrease (increase) in the external excitatory input. Generalizing the notion of paradoxical responses of inhibitory neurons [10], we denote any negative response as paradoxical. We will in particular examine the relation between local connectivity statistics and paradoxical responses.
@@ II Low-rank approximation
While there exists a variety of methods for forming a low-rank approximation, here we use a simple truncated eigen-decomposition. We therefore take the vectors $\bm{n}^{(r)},~{}\bm{m}^{(r)}\in\mathbb{R}^{N}$ to be respectively the $r-$th left and right eigenvectors $\bm{L}_{r}$ and $\bm{R}_{r}$ re-scaled by $\sqrt{N}$:
@@ III.1 General approach
$=\det(\mathbf{Z}-\mathbf{1}\lambda)\frac{1}{\lambda}\det\left(\lambda\mathbf{1}-\frac{1}{N}\mathbf{N}_{0}^{\intercal}(\mathbf{1}-\mathbf{Z}/\lambda)^{-1}\mathbf{M}_{0}\right).$
@@ III.2.1 Definition
Figure 2: Impact of chain motifs on the eigenvalues of the connectivity matrix for fully connected excitatory-inhibitory networks. (a) Eigenspectrum in the complex plane. The spectrum consists of a circular bulk (magnification in inset), within which the eigenvalues are continuously distributed, and isolated outliers. Inhibition dominated networks with independent synapses give rise to a single negative outlier $\lambda_{0}$ (black circle). Progressively increasing the strength $\tau^{c}$ of chain motifs (from light to dark), this negative outlier (yellow to red) decreases from the original $\lambda_{0}$, while an additional positive outlier emerges and increases (light to dark green). The dots show numerically-determined outlying eigenvalues averaged over $30$ networks of $N=1000$ neurons. The inset shows the eigenvalue bulk for a single network realization with $\tau^{c}=0$ (purple) and $\tau^{c}=0.1$ (blue). (b) Dependence of outlying eigenvalues on the strength $\tau^{c}$ of chain motifs and the network size $N$. Numerically obtained eigenvalue outliers from $30$ network realizations (markers) are shown alongside the theoretical predictions (lines) calculated using Eq.(34). The dominant negative outlier is depicted in red, while the emergent positive outlier is denoted in green. The gray area indicates the radius of the eigenvalue bulk in networks with a size of $N=1000$, with dashed lines indicating the theoretical values [30]. All parameter values are given in TABLE 2.
@@ III.2.2 Eigenvalues
In comparison with chain motifs the perturbation term is independent of $N$, so that reciprocal motifs have a much weaker effect on outlying eigenvalues [33]. This difference can be traced back to Eq. (26), where reciprocal motifs appear only in the $N$ diagonal entries of $\mathbf{D}$, while chain motifs determine the $N(N-1)$ off-diagonal elements in $\mathbf{U}_{o}\mathbf{V}_{o}^{\intercal}$, which leads to a factor $N-1$ in Eq.(35). Put otherwise, there are many more possible chain than reciprocal motifs in a circuit.
@@ III.3.4 Weakly connected regime
We next consider a second type of scaling with network size, where the connection probability $c$ and chain motif occurence $\rho^{c}$ decrease with $N$ in such a way that the number $C_{p}$ of incoming synaptic connections for each neuron, the number of motifs $k^{c}_{pq}$ and the synaptic strengths $J$ are fixed independently of $N$. This scaling is analogous to the weakly connected regime in networks without motifs [58]. Here, $C_{p}$ represents the number of excitatory (if $p=E$) or inhibitory (if $p=I$) synapses targeting the same neuron, while $k^{c}_{pq}$ denotes the number of chain motifs originating from neurons in population $q$, passing through neurons in population $p$ and ultimately targeting an individual neuron. These parameters are related to the connection probability $c$ and motif occurrence probability $\rho^{c}$ via
@@ IV Population-averaged responses to uniform inputs
In this section, we use a low-rank approximation of recurrent connectivity to determine how chain motifs affect steady-state responses to external inputs. We focus on population-averaged responses to inputs that are uniform over each population, and specifically examine under which conditions paradoxical responses emerge. We start by outlining the general approach, and then apply it to fully-connected and sparse excitatory-inhibitory networks.
@@ IV.1 Low-rank approximation
IV.2 Fully connected Gaussian excitatory-inhibitory networks
Figure 6: Impact of chain motifs on the response to uniform external inputs in fully-connected Gaussian excitatory-inhibitory networks: mean response $\chi_{E}$ of the excitatory population as a function of the strength $\tau^{c}$ of chain motifs, which is homogeneous across all blocks of the connectivity matrix. The asterisks with error bars show the mean and s.d. of population-averaged responses in 30 realizations of the full connectivity (Eq. (10)). The lines represent the predictions of the low-rank approximation in Eq. (50), where the $a_{x}^{p}$ were obtained numerically by diagonalizing the full connectivity matrix. The dashed line is the prediction obtained from the mean connectivity $\mathbf{J}^{0}$ (Eq. (51)). The colored light lines show the respective contributions of the two unit-rank terms in the low-rank approximation of the response (Eq. (50)). Orange and green denote respectively the modes corresponding to the negative and positive outliers. Inset: outlying eigenvalues as a function of $\tau^{c}$, compared to the bulk of the radius. The positive outlier emerges from the bulk for $\tau^{c}=0.011$. All network parameters are given in TABLE 2.
@@ IV.2 Fully connected Gaussian excitatory-inhibitory networks
IV.2.1 Response to uniform inputs
Figure 7: Impact of chain motifs on the mean response $\chi_{pq}$ of population $p$ to uniform inputs to population $q$, in fully-connected Gaussian excitatory-inhibitory networks. (a) $\chi_{EE}$; (b) $\chi_{IE}$; (c) $\chi_{EI}$; (d) $\chi_{II}$. The asterisks with error bars show the mean and s.d. of population-averaged responses in $30$ realizations of the full connectivity (Eq. 10). The solid lines represent the predictions of the low-rank approximation (Eq. (48)), with colored lines indicating the contributions from the two unit rank terms (orange: negative outlier; green: positive outlier). In Eq. (48), $a_{x}^{p}$ were obtained numerically by diagonalizing the full connectivity matrix. The strength $\tau^{c}$ of chain motifs is homogeneous across all blocks of the connectivity matrix. The mean connectivity parameters were chosen such that $\chi_{II}>0$ for $\tau^{c}=0$, i.e. the network is in the non-paradoxical regime in absence of inputs. All network parameters are given in TABLE 2.
@@ IV.2.1 Response to uniform inputs
While the mean connectivity $\mathbf{J^{0}}$ predicts accurately the mean response of random networks without correlations between synapses, we find that including chain motifs strongly amplifies the response of both excitatory and inhibitory populations (Fig 6). This amplification is well captured by the low-rank approximation, which is in excellent agreement with the activity in the full model as $\tau^{c}$ is increased. The accuracy of the low-rank approximation can be explained by the fact that the eigenmodes which correspond to the bulk eigenvalues that were not included in the low-rank description have zero mean components, and therefore do not contribute to the population-averaged response (see insets in Fig 5 (b,c)).
Within the low-rank approximation, we can distinguish the contributions to the response of the two unit-rank terms (Fig 6 colored lines). This decomposition reveals that the unit-rank term associated with the negative outlying eigenvalue has minimal influence on the amplification induced by chain motifs. Conversely, the amplification is due to the second term that corresponds to the newly emerging eigenmode with a positive outlier. The amplification is therefore specific to the outlier induced by chain motifs. In contrast, networks featuring only reciprocal motifs do not give rise to this outlier, and therefore, there is no noticeable amplification as $\tau^{r}$ increases (Fig D.2(a, b)).
IV.2.2 Paradoxical responses
We next use the low-rank approximation to examine the mean responses $\chi^{pq}$ of population $p$ to uniform inputs to population $q$. Fig 7 shows that chain motifs strongly modulate all components of this response matrix. Focusing on the diagonal entries we denote the response of population $p$ as paradoxical if $\chi_{pp}<0$, i.e. if it responds in a direction opposite to the input it receives. Fig 7 (d) shows that increasing the strength of chain motifs changes the sign of $\chi_{II}$ from positive to negative, thereby switching the response of inhibitory neurons from non-paradoxical in the absence of chain motifs to paradoxical as $\tau^{c}$ is increased.

### v3

@@ Introduction
These seminal theoretical results on excitatory-inhibitory networks however are derived via a key simplifying assumption. In standard models, the strength of the synaptic coupling between any two neurons depends on their types, but is otherwise assumed to be an independent random variable uncorrelated across synapses. This assumption is typically used to reduce a full network to a simpler circuit model that describes how the mean activities of different populations interact through averaged synaptic weights (Fig 1(a,b). Recent synaptic-resolution experimental datasets from various species and brain areas have however revealed the prevalence of non-trivial connectivity patterns [22, 23, 24, 25, 26, 27, 28, 29, 30]. In particular, a recently released dataset from mice and humans [23, 30] reported the prominence of second-order motifs - specific correlations between pairs of synapses, such as reciprocal, chain, convergent, and divergent motifs (Fig 1(c)) - that go well beyond what is expected from independent connections, and highlights the need for a theoretical understanding of the effects of such patterns. The presence of reciprocal connectivity motifs has been long recognized and examined within network models [29, 31, 30, 32, 33]. Other studies have argued for the importance of several types of synaptic motifs working together to determine statistical properties of network activity, such as average synchrony or correlation among neurons and network-wide dimensionality [34, 35, 32, 30, 36, 37, 38, 39, 40]. In a general theoretical analysis [35], chain motifs – corresponding to a pattern where neurons with stronger inputs also have stronger outputs – were found to have a dominant role in determining population-averaged responses of networks to their inputs. Nevertheless, the impact of chain motifs on excitation-inhibition balance – including central issues of stability and paradoxical responses – as well as the general interplay among these motifs and structures established by cell-type specific connectivity, remain open and intriguing questions.
@@
Figure 1: Schematic of the multi-population network model. (a) The network consists of $P$ populations, each represented by a different color. This population structure defines the statistics of synaptic connectivity. (b) The corresponding connectivity matrix consists of $P^{2}$ blocks. Synaptic weights within each block share identical statistical properties. (c) Pairs of synapses that share a neuron can form four different types of second-order motifs. The prevalence of each motif with respect to chance is quantified by a corresponding pair-wise correlation coefficient. (d) A low-rank approximation of the connectivity matrix can integrate both the population structure and the pair-wise motif statistics.
@@
In this study, we examine how the interaction between population structure and synaptic motifs influences recurrent dynamics and the presence of paradoxical responses. To this end, we expand a previously introduced theoretical framework that allows us to reduce large networks of multiple populations to a low-dimensional description that incorporates connectivity statistics beyond the mean via a low-rank approximation [33]. Applying this theory to excitatory-inhibitory networks with chain motifs, we demonstrate that these connectivity patterns significantly impact the eigen-spectrum of the connectivity matrix, and thereby the overall recurrent dynamics. Specifically, an over-representation of chain motifs creates strong positive feedback even in inhibition-dominated networks, and therefore leads to a potential instability that requires revisiting the classical conditions of excitation-inhibition balance. Moreover, we found that chain motifs strongly influence the responses of different populations to external inputs, and can control whether the responses are paradoxical or not. We show that the paradoxical responses can be equivalently predicted from two approaches: (i) a low-rank approximation based on the outliers in the eigenspectrum; (ii) an effective connectivity matrix which combines average synaptic weights and the strength of chain motifs. Altogether, our findings highlight the intricacy of the relationship between the responses to inputs and the underlying connectivity, and in particular sound a note of caution for interpreting results of experiments in terms of only average connectivity strengths among E and I cells.
@@ I.1 Network connectivity
Following a common approach [29, 41, 34, 36, 37, 38, 39, 32], we begin by characterizing connectivity through the marginal distributions $\mathcal{P}(J_{ij}=J)$ of individual synaptic weights, and progressively incorporate higher-order correlations. In this study, we consider the first and second orders, that is, the marginal distribution of individual synaptic weights and the pair-wise correlations. We denote as motifs correlations between weights of pairs of synapses that share a neuron (Fig 1(c)). We specifically focus on chain motifs corresponding to correlations between synapses $J_{ij}$ and $J_{jk}$ which share the intermediate neuron $j$, but $i\neq k$. The strength of chain motifs can be quantified by the correlation coefficient
@@
This measure of motif strength takes the form of a Pearson correlation coefficient, representing the prevalence of correlations between a pair of synaptic weights compared to the independent case. Here, and throughout this study, square brackets denote the average over the full connectivity distribution.
@@
Pairs of synapses sharing a neuron can form two additional types of correlations corresponding to convergent and divergent motifs. We show in Appendix. A that, in the limit of large networks, they do not contribute to the dynamics we study here.
@@ I.2 Network dynamics
Our aim is to determine how external inputs to different populations affect the steady-state activity. To this end, we will examine the response matrix $\bm{\chi}$, where the entry $\chi_{ij}$ represents the change of the steady-state activity of neuron $i$ resulting from a change in the input current to neuron $j$:
@@
$\bm{\chi}=(\mathbf{1}-\mathbf{J})^{-1}.$ (10)
The diagonal entries of the response matrix, denoted as $\chi_{ii}$, represent the activity change of neuron $i$ in response to the direct external input they receive. When this term is negative, this indicates that the neuron increases (decreases) its firing rate with a decrease (increase) in the external excitatory input. Generalizing the notion of paradoxical responses of inhibitory neurons [10], we denote any negative diagonal element of the response function as a paradoxical response.
To examine the relationship between local connectivity statistics and paradoxical responses to external inputs, we compare two approaches: (i) responses obtained from a low-rank approximation based on the dominant eigenmodes of the random connectivity matrix $\mathbf{J}$ [33]; (ii) a direct approximation of the average response function in which the random connectivity matrix $\mathbf{J}$ is replaced by a deterministic matrix $\mathbf{J}^{eff}$ that combines the first and second order statistics [35].
@@ II Low-rank approximation
While there exists a variety of methods for forming a low-rank approximation, here we use a simple truncated eigen-decomposition, which assumes that the connectivity matrix $\mathbf{J}$ is diagonalizable. Specficially, we take the vectors $\bm{n}^{(r)},~{}\bm{m}^{(r)}\in\mathbb{R}^{N}$ to be respectively the $r-$th left and right eigenvectors $\bm{L}_{r}$ and $\bm{R}_{r}$ re-scaled by $\sqrt{N}$:
@@ III.1 General approach
$=\det(\mathbf{Z}-\mathbf{1}\lambda)\frac{1}{\lambda^{N}}\det\left(\lambda\mathbf{1}-\frac{1}{N}\mathbf{N}_{0}^{\intercal}(\mathbf{1}-\mathbf{Z}/\lambda)^{-1}\mathbf{M}_{0}\right).$
@@ III.2.1 Definition
Figure 2: Impact of chain motifs on the eigenvalues of the connectivity matrix for fully connected excitatory-inhibitory networks. (a) Eigenspectrum in the complex plane. The spectrum consists of a circular bulk (magnification in inset), within which the eigenvalues are continuously distributed, and isolated outliers. Inhibition dominated networks with independent synapses give rise to a single negative outlier $\lambda_{0}$ (black circle). Progressively increasing the strength $\tau^{c}$ of chain motifs (from light to dark), this negative outlier (yellow to red) decreases from the original $\lambda_{0}$, while an additional positive outlier emerges and increases (light to dark green). The dots show numerically-determined outlying eigenvalues averaged over $30$ networks of $N=1000$ neurons. The inset shows the eigenvalue bulk for a single network realization with $\tau^{c}=0$ (purple) and $\tau^{c}=0.1$ (blue). (b) Dependence of outlying eigenvalues on the strength $\tau^{c}$ of chain motifs and the network size $N$. Numerically obtained eigenvalue outliers from $30$ network realizations (markers) are shown alongside the theoretical predictions (lines) calculated using Eq.(34). The dominant negative outlier is depicted in red, while the emergent positive outlier is denoted in green. The gray area indicates the radius of the eigenvalue bulk in networks with a size of $N=1000$, with dashed lines indicating the theoretical values [30]. As a control, we fix $\lambda_{0}$ across networks with different $N$, by scaling the mean synaptic weight $J^{0}_{pq}$ as $1/N$. All parameter values are given in TABLE 2.
@@ III.2.2 Eigenvalues
Compared to chain motifs, the perturbation term is independent of $N$, so that reciprocal motifs have a much weaker effect on outlying eigenvalues [33]. This difference can be traced back to Eq. (26), where reciprocal motifs appear only in the $N$ diagonal entries of $[\mathbf{Z}^{2}]$, while chain motifs determine the $N(N-1)$ off-diagonal elements, which leads to a factor $N-1$ in Eq.(35). Convergent and divergent motifs have even weaker effects on outlying eigenvalues (Fig. E.1), as they do not appear in $[\mathbf{Z}^{2}]$, and enter only the averages of higher powers of $\mathbf{Z}$ at the sub-dominant order in $1/N$ (Appendix A).
@@ III.3.4 Weakly connected regime
We next consider a second type of scaling with network size, where the connection probability $c$ and chain motif occurrence $\rho^{c}$ decrease with $N$ in such a way that the number $C_{p}$ of incoming synaptic connections for each neuron, the number of motifs $k^{c}_{pq}$ and the synaptic strengths $J$ are fixed independently of $N$. This scaling is analogous to the weakly connected regime in networks without motifs [58]. Here, $C_{p}$ represents the number of excitatory (if $p=E$) or inhibitory (if $p=I$) synapses targeting the same neuron, while $k^{c}_{pq}$ denotes the number of chain motifs originating from neurons in population $q$, passing through neurons in population $p$ and ultimately targeting an individual neuron. These parameters are related to the connection probability $c$ and motif occurrence probability $\rho^{c}$ via
@@ IV Population-averaged responses to uniform inputs
In this section, we examine how chain motifs affect steady-state responses to external inputs. We focus on population-averaged responses to inputs that are uniform over each population, and specifically examine under which conditions paradoxical responses emerge. To this end, we compare two methods: (i) predicting responses from a low-rank approximation of the recurrent connectivity matrix; (ii) predicting responses from an effective deterministic matrix $\mathbf{J}^{eff}$. We start by outlining the general approach, and then apply it to fully-connected and sparse excitatory-inhibitory networks.
@@ IV.1 Low-rank approximation
IV.2 Effective connectivity approximation
An alternative approach is to directly average the response function in Eq. (10) over realizations of the random connectivity [35]. We show in Appendix D that, at leading order in $N$, the average response is equivalent to the response generated by the effective connectivity matrix
$\mathbf{J}^{eff}=\mathbf{J}^{0}+\left[\mathbf{Z}^{2}\right].$ (50)
This effective connectivity is a deterministic matrix consisting in general of $P\times P$ blocks. Within each block, all entries are equal, and given by a combination of the mean synaptic strength between two populations and the strength of chain and reciprocal motifs that determine the entries of $\left[\mathbf{Z}^{2}\right]$ (Eq. (25)). $\mathbf{J}^{eff}$ is therefore a low-rank matrix, of rank at most $P^{2}$, and equivalent to a $P\times P$ matrix. The response predicted by $\mathbf{J}^{eff}$ can then be analyzed using methods developed for deterministic connectivity matrices [10, 15].
IV.3 Fully connected Gaussian excitatory-inhibitory networks
Figure 6: Impact of chain motifs on the response to uniform external inputs in fully-connected Gaussian excitatory-inhibitory networks: mean response $\chi_{E}$ of the excitatory population as a function of the strength $\tau^{c}$ of chain motifs, which is homogeneous across all blocks of the connectivity matrix. The asterisks with error bars show the mean and s.d. of population-averaged responses in 30 realizations of the full connectivity (Eq. (10)). The lines represent the predictions of the low-rank approximation in Eq. (51), where the $a_{x}^{p}$ were obtained numerically by diagonalizing the full connectivity matrix. The gray dashed line is the prediction obtained from the mean connectivity $\mathbf{J}^{0}$ (Eq. (52)), the orange dashed line is the prediction obtained from the effective connectivity $\mathbf{J}^{eff}$. The colored light lines show the respective contributions of the two unit-rank terms in the low-rank approximation of the response (Eq. (51)). Red and green denote respectively the modes corresponding to the negative and positive outliers. Inset: outlying eigenvalues as a function of $\tau^{c}$, compared to the bulk of the radius. The positive outlier emerges from the bulk for $\tau^{c}=0.011$. All network parameters are given in TABLE 2.
@@ IV.3 Fully connected Gaussian excitatory-inhibitory networks
IV.3.1 Response to uniform inputs
Figure 7: Impact of chain motifs on the mean response $\chi_{pq}$ of population $p$ to uniform inputs to population $q$, in fully-connected Gaussian excitatory-inhibitory networks. (a) $\chi_{EE}$; (b) $\chi_{IE}$; (c) $\chi_{EI}$; (d) $\chi_{II}$. The asterisks with error bars show the mean and s.d. of population-averaged responses in $30$ realizations of the full connectivity (Eq. (10)). The orange dashed lines represent the predictions of the effective connectivity $\mathbf{J}^{eff}$, the black triangle in (d) is the prediction of the transition point using the stability of the E-E subnetwork. The solid lines represent the predictions of the low-rank approximation (Eq. (48)), with colored lines indicating the contributions from the two unit rank terms (red: negative outlier; green: positive outlier). In Eq. (48), $a_{x}^{p}$ were obtained numerically by diagonalizing the full connectivity matrix. The strength $\tau^{c}$ of chain motifs is homogeneous across all blocks of the connectivity matrix. The mean connectivity parameters were chosen such that $\chi_{II}>0$ for $\tau^{c}=0$, i.e. the network is in the non-paradoxical regime in absence of inputs. All network parameters are given in TABLE 2.
@@ IV.3.1 Response to uniform inputs
While the mean connectivity $\mathbf{J^{0}}$ predicts accurately the mean response of random networks without correlations between synapses, we find that including chain motifs strongly amplifies the response of both excitatory and inhibitory populations (Fig 6). This amplification is well captured both by the low-rank approximation and the effective connectivity $\mathbf{J}^{eff}$. The accuracy of the low-rank approximation can be explained by the fact that the eigenmodes which correspond to the bulk eigenvalues that were not included in the low-rank description have zero mean components, and therefore do not contribute to the population-averaged response (see insets in Fig 5 (b,c)).
Within the low-rank approximation, we can distinguish the contributions to the response of the two unit-rank terms (Fig 6 colored lines). This decomposition reveals that the unit-rank term associated with the negative outlying eigenvalue has minimal influence on the amplification induced by chain motifs. Conversely, the amplification is due to the second term that corresponds to the newly emerging eigenmode with a positive outlier. The amplification is therefore specific to the outlier induced by chain motifs. In contrast, networks featuring only reciprocal motifs do not give rise to this outlier, and therefore, there is no noticeable amplification as $\tau^{r}$ increases (Fig E.2(a, b)).
IV.3.2 Paradoxical responses
We next use the low-rank approximation to examine the mean responses $\chi_{pq}$ of population $p$ to uniform inputs to population $q$. Fig 7 shows that chain motifs strongly modulate all components of this response matrix. Focusing on the diagonal entries we denote the response of population $p$ as paradoxical if $\chi_{pp}<0$, i.e. if it responds in a direction opposite to the input it receives. Fig 7 (d) shows that increasing the strength of chain motifs changes the sign of $\chi_{II}$ from positive to negative, thereby switching the response of inhibitory neurons from non-paradoxical in the absence of chain motifs to paradoxical as $\tau^{c}$ is increased.